OpenStudy (eriyalovesu):

Which of the following options results in a graph that shows exponential growth? f(x) = 0.4(0.2)x f(x) = 4(0.98)x f(x) = 0.7(5)x f(x) = 5(4)−x

OpenStudy (mathmale):

If you're studying exponential functions, it'd be worth your while to express such functions properly. Use the symbol ^ to indicate exponentiation: x squared would be x^2. By f(x) = 4(0.98)x you are apparently indicating exponentiation: f(x) = 4(0.98)^x Graphs of exponential functions show either growth or decay. In this case you're interested in identifying which given function has a graph that resembles the following in shape: 

OpenStudy (mathmale):

1. Please re-write your four functions to indicate exponentiation properly. 2. Graph all four functions. 3. Identify the one that looks most like the illustration, above.

OpenStudy (sepeario):

lol I think they're copied and pasted, that's why they don't have the superscript.